OpenStudy (anonymous):

Use the boxplot values to determine the shape of the distribution. minimum = 15.8 Q1 = 23.6 median = 29.2 Q3 = 34.1 maximum = 54.3 Question 10 options: left-skewed right-skewed approximately normal

OpenStudy (amistre64):

where is the median as far as the range goes?

OpenStudy (amistre64):

(54.3+15.8)/2 = 35.05 the median is to the left of the mean; so it would be right skewed if i see it corectly

OpenStudy (amistre64):

but when you take the outliers off; your pretty much normalized then

OpenStudy (amistre64):

so i have to wonder if i recall these correctly

OpenStudy (anonymous):

where do you see outliers

OpenStudy (amistre64):

min and max are outliers; which is why you dont use them in the interquartile range

OpenStudy (amistre64):

but the position of the box itself is the clue and it is off the middle

OpenStudy (amistre64):

its heavy to the left so it skews right

OpenStudy (anonymous):

i agree skewed right

OpenStudy (amistre64):



OpenStudy (anonymous):

i thought one measure of outliers using box plots was if the values was more than 1.5*IQR's from Q1 or Q3, or am i remembering incorrectly

OpenStudy (amistre64):

dunno, you seem to be remembering more than me so id say go for it :)

OpenStudy (anonymous):

IQR=15.75 Q3+1.5 IQR=34.1+1.5*15.75=57.727

OpenStudy (anonymous):

and the max is 54.3, i.e., not an outlier; btw there is another measure for an "extreme" outlier but I do not remember the scale factor for the IQR